Lactic acid, C3H6O3, is built in human muscles during anaerobic exercises. If the initial concentration of lactic acid is 0.12 mol/L and the pH is 2.39, what is the Ka for the lactic acid?




1
Expert's answer
2021-06-25T06:42:23-0400

The pH is telling us the concentration of the H3O+(aq) and C3H5O3(aq)

pH =–log [H3O+(aq)]

[H3O+] =10– pH = 10–2.39 = 4.1 × 10–3 mol/L

Ka = ([H3O+][C3H5O3]) / [C3H6O3] = ((4.1×10−3)(4.1×10−3)) / 0.116 = 1.45 × 10−4

Therefore, Ka is 1.45 × 10−4.